### v2

# Strong Gravitational Lensing Effects of the Rotating Short-Haired Black Hole and Constraints from EHT Observations

Abstract
For the short hairs that have a significant impact only near the event horizon, studying their strong gravitational lensing effects is of great significance for revealing the properties of these hairs. In this study, we systematically investigated the strong gravitational lensing effects in the rotating short-haired black hole and constrained its hair parameter $Q_{m}$. Specifically, $Q_{m}$ causes the event horizon radius, photon - orbit radius, and impact parameter to be lower than those of the Kerr black hole. Regarding the lensing coefficients $\bar{a}$ and $\bar{b}$, as the spin parameter $a$ increases, $\bar{a}$ shows an increasing trend, while $\bar{b}$ shows a decreasing trend. In the observational simulations of M87* and Sgr A*, the angular position and angular distance of the relativistic image increase with the increase of $a$, while the magnification of the image shows an opposite trend. The existence of $Q_{m}$ only intensifies these trends. More importantly, the rotating short-haired black hole exhibits a significant difference in time delay compared to other black hole models. Especially in the simulation of M87*, the time delay deviation between the rotating short-haired black hole and the Kerr black hole, as well as the Kerr-Newman black hole, can reach dozens of hours. Through a comparative analysis with the observational data from the EHT, we effectively constrain the parameter space of the rotating short-haired black hole. The results indicate that this model has potential application prospects in explaining cosmic black hole phenomena and provides a possible theoretical basis for differentiating between different black hole models.
Keywords
Strong Gravitational Lensing; Rotating Short-haired Black Hole;Time Delay
preprint: APS/123-QED
I Introduction
The general theory of relativity predicts the existence of black holes(BH), but finding evidence of BH in the real universe is crucial for validating this theory. This has greatly stimulated the interest of physicists. In September 2015, the LIGO detector first captured gravitational wave signals produced by the merger of two BHs, providing strong evidence for the success of general relativity Abbott et al. (2016). In the real universe, most of the observed astrophysical BHs are rotating and fit well with the Kerr metric Teukolsky (2015); Tao et al. (2023). However, this does not rule out the possibility of a range of Kerr-like BHs Azreg-Aïnou (2014); Ghosh (2015). In fact, as Xu et al. stated in the literature Xu et al. (2021), regarding the Kerr BH as the astrophysical BH in the real universe is not entirely applicable. This is because BH in the real universe do not exist in isolation but are coupled with dark matter or other forms of fields, making the real universe more complex. Therefore, accurately describing astrophysical BH using standard Kerr BH may present certain challenges.
If a type of Kerr-like BH solution that considers coupling with anisotropic matter could be found to describe astrophysical BH, it might be more meaningful. The short-haired BH solution provided in the literature Brown and Husain (1997) better fits the real universe environment because their short-hairy BH solution considers BH solutions coupled with anisotropic matter. Their static spherically symmetric BH solution violates the no-hair theorem. The no-hair theorem states Hawking (1972); Israel (1967) that within the framework of general relativity, the properties of a BH can be completely described by three parameters: mass $M$, charge $Q$, and angular momentum $J$. However, once Einstein’s gravity is coupled with other matter, BH solutions that violate the no-hair theorem may emerge, such as Lee et al. (2019); Greene et al. (1993); Bizon (1990); Ovalle et al. (2021). Therefore, studying the properties of short-hairy BH solutions Brown and Husain (1997) would be very interesting. However, the BH solution they provided is static and spherically symmetric. It would be more valuable to extend this solution to the rotating and axisymmetric case. This is because, on the one hand, most astrophysical BHs are rotating, so a rotating solution can better match astrophysical BHs. On the other hand, jet phenomena are common around rotating BHs, and observing these jets helps us study and understand the physical behavior near the event horizon, especially for this type of short hair that has a significant impact near the event horizon. Based on this objective, Tang and Xu in 2022 used the Newman-Janis (NJ) algorithm to extend the short-hairy BH to the case of rotating short-hairy BH, and they also studied the impact of short-hair parameter on the BH shadow Tang and Xu (2022). Studying other properties of the rotating short-hairy BH is also very meaningful. For example, exploring their physical behavior under strong gravitational lensing can help us further understand the characteristics of rotating short-hairy BH and provide a window to test the no-hair theorem.
Gravitational lensing is an important tool for studying black holes (or other massive objects) and their surrounding environments. It has been widely applied not only in understanding the structure of spacetime (e.g., Liebes (1964); Mellier (1999); Guzik et al. (2010); Schmidt (2008)), but also plays a crucial role in the search for dark matter (see Massey et al. (2010); Vegetti et al. (2023); Diaz Rivero and Dvorkin (2020); Şengül and Dvorkin (2022); Fairbairn et al. (2023), etc.). Meanwhile, it is an effective means to test gravitational theories (see Zhao et al. (2024); Bekenstein and Sanders (1994); Islam et al. (2020); Kumar et al. (2020); Chen and Jing (2009), etc.). In a strong gravitational field, such as near a BH, when light rays approach the BH (especially near the photon sphere), the light rays will be strongly bent, and may even orbit the black hole one or more times, that is, the deflection angle of the light rays exceeds $2\pi$. In this case, shadows, photon rings, and relativistic images will appear Gralla et al. (2019); Bozza (2010); Bozza et al. (2001); Bozza (2003). Regarding strong gravitational lensing, Virbhadra and Ellis analyzed the strong gravitational effects caused by a Schwarzschild BH through numerical simulations Virbhadra and Ellis (2000). Subsequently, Bozza Bozza et al. (2001); Bozza (2002, 2010) and Tsukamoto Tsukamoto (2017), among others, generalized it to a general static spherically symmetric spacetime. Of course, other static spherically symmetric BHs have also been studied accordingly (e.g., Man and Cheng (2015); Wei et al. (2015); Eiroa and Sendra (2013, 2012); Eiroa (2006, 2005); Zhao and Xie (2017); Zhao et al. (2024)). In addition, rotating axisymmetric black holes have also received extensive attention, such as Bozza (2008a, 2010, 2003); Wei et al. (2012); Islam and Ghosh (2021); Islam et al. (2021); Hsiao et al. (2020).
In fact, in 2019, the EHT successfully captured images of the supermassive BH $M87^{*}$ Akiyama et al. (2019a), and in 2022, it captured images of the BH at the center of the Milky Way, $SgrA^{*}$ Akiyama et al. (2022a). These breakthrough observations have triggered an upsurge of research on strong gravitational lensing. The EHT’s observations provide a new laboratory for studying the properties of black holes, such as the black hole’s shadow, the jet phenomenon of the accretion disk, the event horizon, and other properties that previously could only be theoretically calculated but can now be verified through observation. Naturally, studying the strong gravitational lensing effects around BH is also of great significance, as it can magnify and distort the images of background celestial bodies, providing a unique method to detect BHs and their surrounding material distribution. Therefore, studying strong gravitational lensing in BHs that are more akin to those in the real universe holds greater physical value. The environment considered for the rotating short-haired BH Tang and Xu (2022) is more similar to the real universe (anisotropic matter). Studying strong gravitational lensing in such BH can help us test the no-hair theorem, distinguish Kerr BH, and further understand the properties of short-hair near the event horizon. For this kind of hair, which has significant effects near the event horizon but is difficult to observe for distant observers, observing the behavior of light reaching near the event horizon will help us better understand the properties of this short-hair.
The paper is organized as follows: In Section II, the short-haired BH and its event horizon information are briefly introduced. In Section III, we calculate the lensing coefficients of the rotating short-haired BH respectively, and explore the influence of the short-hair parameter $Q_{m}$ on its lensing. In Section IV, considering the rotating short-haired BH as a candidate for the supermassive BHs M87* and Sgr A*, we discuss the lensing observational effects and time delay of the rotating short-haired BH. In Section V, we explore the use of observational data to systematically constrain the parameter space of the rotating short-haired BH. In Section VI, necessary discussions are carried out. In this paper, we adopt the natural unit system, i.e., $c=G=\hbar=1$.
II Rotating Short-Haired Black Hole
In the framework of classical general relativity, the no-hair theorem states that the properties of a BH can be fully described by just three parameters: mass $M$, charge $Q$, and angular momentum $J$ Hawking (1972); Israel (1967). However, research indicates that the no-hair theorem might not be universal, and numerous studies have sought counterexamples. In the 1990s, Piotr Bizon was the first to discover a BH solution with ”hair” through numerical analysis, challenging the absoluteness of the no-hair theorem Bizon (1990). Ovalle et al. constructed a spherically symmetric BH solution with hair via the gravitational decoupling method Ovalle et al. (2021), and this was later extended to the rotating case Contreras et al. (2021). Additionally, other researchers have proposed BH solutions with hair based on scalar fields or other theoretical frameworks, such as Caramês (2023); Bakopoulos and Nakas (2023); Zhang (2022); Brihaye and Ducobu (2016). In the literature Brown and Husain (1997), the authors obtained short-haired BH solutions by coupling Einstein’s gravity with anisotropic matter. These solutions can present de Sitter and Reissner-Nordström BHs in some cases, and in others, they yield short-haired BH solution Brown and Husain (1997). The corresponding metric is given by
$f(r)=1-\frac{2M}{r}+\frac{Q_{m}^{2k}}{r^{2k}}.$ (1)
Among them, $Q_{m}$ is the hair strength parameter, which may represent certain quantum effect related hairs and has a significant impact on the structure near the event horizon.
Considering that actual BHs in the universe usually have rotational characteristics, it is particularly important to generalize such short-haired BH solutions to the rotating case. Against this background, Tang and Xu generalized the static spherically symmetric short-haired BH metric to the rotating case through the Newman-Janis (NJ) algorithm, obtained the rotating short-haired BH solution, and analyzed the influence of the short- hair parameter on the BH shadow Tang and Xu (2022). This generalization makes it possible to study the physical properties of short-haired BHs in a rotating background. Since the spacetime of a rotating short-hairedBH is closer to that of the real universe, the matter distribution within it is non-vacuum and anisotropic. This model provides a basis for further exploring the properties of rotating short-haired BHs. The metric form of the rotating short-haired BH is Tang and Xu (2022)
$ds^{2}$ $=-\left(1-\frac{2Mr-\frac{Q_{m}^{2k}}{r^{2k-2}}}{\rho^{2}}\right)dt^{2}+\frac{\rho^{2}}{\Delta}dr^{2}$
$-\frac{2a\sin^{2}\theta\left(2Mr-\frac{Q_{m}^{2k}}{r^{2k-2}}\right)}{\rho^{2}}dtd\phi+\rho^{2}d\theta^{2}$
$+\frac{\Sigma\sin^{2}\theta}{\rho^{2}}d\phi^{2},$ (2)
where
$\rho^{2}=r^{2}+a^{2}\cos^{2}\theta,$ (3)
$\Sigma=(r^{2}+a^{2})^{2}-a^{2}\Delta,$ (4)
$\Delta=r^{2}-2Mr+\frac{Q_{m}^{2k}}{r^{2k-2}}+a^{2}.$ (5)
Here, $a$ denotes the black hole’s spin parameter, and $Q_{m}$ represents the hair strength parameter. As analyzed in the original literature Brown and Husain (1997), the short-haired BH satisfies the Weak Energy Condition (WEC). Thus, the value of the parameter $k$ must meet the requirement $2k-1\geq 0$, meaning $k\geq\frac{1}{2}$. Moreover, the trace of the energy momentum tensor $T=2\rho(k-1)>0$ (this holds when $k>1$), indicating that the short-haired BH does not violate the no ”short-hair” theorem Nunez et al. (1996); Acharya and Sarkar (2024). When $k=1$, the metric (2) simplifies to the classical Kerr-Newman BH. When $k>1$, the metric (2) depicts a short-haired BH. In the discussion of this paper, we will concentrate on analyzing the BHs corresponding to $k=1$ and $k=\frac{3}{2}$, with particular emphasis on the case of $k=\frac{3}{2}$. The reason is that as the value of $k$ increases, the influence of the hair strength parameter $Q_{m}$ gradually diminishes, and the properties of the short-haired BH gradually approximate those of the Kerr BH, making it difficult to distinguish between them. Therefore, choosing $k=\frac{3}{2}$ as the main focus of our study can more clearly expose the characteristics of the short-haired BH and its differences from classical BHs (such as the Kerr-Newman BH with $k=1$ and the Kerr BH with $Q_{m}=0$).
To facilitate subsequent analysis, we’ve performed a non-dimensionalization of the metric (2), using $2M$ as the unit to non-dimensionalize physical quantities (e.g., $r\rightarrow 2Mx$). In this framework, for the parameter models under discussion, the event horizon of the rotating short-haired BH can be determined by the condition $g^{rr}=0$, which is equivalent to $\Delta=0$. As depicted in Figure 1, the distribution characteristics of the event horizons of rotating short-haired BHs under various conditions are presented. In the figure, the black curve represents the event horizon of the Kerr BH. It can be observed that for different values of the parameter $k$, the presence of the hair parameter $Q_{m}$ significantly impacts the position of the event horizon. Compared to the Kerr BH, the event horizon of the rotating short-haired BH is always smaller. This indicates that the introduction of the hair parameter $Q_{m}$ not only alters the spacetime structure of the BH but also enables in-depth study of the physical properties of the hair parameter within a strong gravitational field environment. Consequently, it offers a new theoretical perspective for understanding the characteristics of short-haired BHs.
[…]
III Strong Gravitational Lensing in the Spacetime of the Rotating Short-hairy Black Hole
For the metric of a rotating BH, to facilitate the calculations, we confine the light rays to the equatorial plane, i.e., $\theta=\pi/2$. The spacetime line element can be written as
$ds^{2}=-A(x)dt^{2}+B(x)dx^{2}+C(x)d\phi^{2}-D(x)dtd\phi.$ (6)
By corresponding the metric of the rotating short-haired BH (2) with equation (6), we can obtain
$A(x)$ $=\left(1-\frac{x-\frac{Q_{m}^{2k}}{x^{2k-2}}}{x^{2}}\right),$ (7)
$B(x)$ $=\frac{x^{2}}{\Delta},$ (8)
$C(x)$ $=\frac{\Sigma}{x^{2}},$ (9)
$D(x)$ $=\frac{2a\left(x-\frac{Q_{m}^{2k}}{x^{2k-2}}\right)}{x^{2}}.$ (10)
For the light rays confined to the equatorial plane, the corresponding Lagrangian is
$L=\frac{1}{2}g_{\mu\nu}\frac{dx^{\mu}}{d\lambda}\frac{dx^{\nu}}{d\lambda},$ (11)
where $\lambda$ is the affine parameter. Due to the axisymmetry of the rotating short-hairy BH, there are two conserved quantities: energy and angular momentum, given by
$E=\frac{\partial L}{\partial\dot{t}}=-g_{tt}\dot{t}-g_{t\phi}\dot{\phi},$ (12)
$L=-\frac{\partial L}{\partial\dot{\phi}}=g_{t\phi}\dot{t}+g_{\phi\phi}\dot{\phi},$ (13)
where the overdot represents differentiation with respect to the affine parameter, i.e., $d/d\lambda$. When choosing an appropriate affine parameter such that $E=1$, the equations of motion for light can be written as
$\dot{t}=\frac{4C-2DL}{4AC+D^{2}},$ (14)
$\dot{\phi}=\frac{2D+4AL}{4AC+D^{2}},$ (15)
$\dot{x}=\pm 2\sqrt{\frac{C-DL-AL^{2}}{B(4AC+D^{2})}}.$ (16)
The effective potential can be defined through equation (16) as
$V_{\text{eff}}(x)=-\dot{x}^{2}=-4\frac{C-DL-AL^{2}}{B(4AC+D^{2})}.$ (17)
By using the impact parameter $u=L/E=L$ to replace the angular momentum $L$, equation (17) can be written as
$V_{\text{eff}}(x)=-4\frac{C-Du-Au^{2}}{B(4AC+D^{2})}.$ (18)
In the above expressions, $A$, $B$, $C$, and $D$ are all functions of $x$.
According to the effective potential, when light rays from infinity approach a rotating short-haired BH, the light rays will be deflected near a specific radius $x_{0}$. Some of the deflected light rays escape to infinity and can be detected by distant observers. When light rays approach the specific orbits of the BH, that is, the photon orbits, these orbits are highly unstable. A slight perturbation can cause the light rays to either be gravitationally captured and fall into the black hole or escape to infinity. The existence of photon orbits determines the boundary of the BH shadow and provides the conditions for the formation of the lensing effect.
The condition for an unstable Photon orbit is given by Harko et al. (2009)
$V_{\text{eff}}(x)=0,\;\left.\frac{dV_{\text{eff}}(x)}{dx}\right|_{x_{m}}=0,\;\left.\frac{d^{2}V_{\text{eff}}(x)}{dx^{2}}\right|_{x_{m}}<0.$ (19)
Combining equations (18) and (19), the orbit equation for the unstable photon orbit can be derived as
$A(x)C^{\prime}(x)-A^{\prime}(x)C(x)+u\left(A^{\prime}(x)D(x)-A(x)D^{\prime}(x)\right)=0.$ (20)
Generally, the above equation has multiple solutions. Only the largest one is outside the event horizon radius, so it is defined as the radius of the unstable photon orbit. This can be clearly seen from the graph of the effective potential. As shown in Figure 2, when the impact parameter reaches the critical value (the green dashed line), the radius of the corresponding unstable photon orbit is marked as $x=x_{m}$. In the following discussion, we’ll use $x_{m}$ to represent the radius of the unstable photon orbit. As presented in Figure 3, because of the hair parameter $Q_{m}$, the radius $x_{m}$ of the unstable photon orbit differs significantly among various models. As the spin parameter increases, $x_{m}$ in different models gradually drops. Besides, the $x_{m}$ of the rotating short-haired BH is smaller than that of the Kerr BH (for the Kerr black hole, $Q_{m}=0$).
Figure 2: The effective potential functions for specific values under different models and the positions of the unstable photon orbits. The intersection point of the green dashed line and the black dashed line represents the position of the unstable photon orbit. Here, we take $Q_{m}=0.3$ and $a=0.2$. From left to right, they correspond to the black - hole models with $k=1$ and $k=1.5$ respectively. We plot the effective potential under different hair parameters.
Figure 3: The variation of the radius of the unstable photon sphere in different BH models with the spin parameter $a$ and the short-hair parameter $Q_{m}$. From left to right, it corresponds to the Kerr-Newman BH with $k=1$ and the rotating short-haired BH with $k=1.5$ respectively.
When a light ray incident from infinity with a certain impact parameter $u$ reaches the vicinity of the BH at $x_{0}$. At this position, the radial velocity of the light ray is zero, while the angular velocity is non-zero. At this time, the light ray is symmetrically deflected to infinity. Since the radial velocity of the light ray is zero at the shortest distance $x_{0}$, the corresponding effective potential reaches its extreme value and satisfies the condition $V_{eff}(x)=0$ (as shown in Figure 2). Through this condition, the relationship between the impact parameter $u$ and the shortest distance $x_{0}$ can be deduced from equation(18) as
$L$ $=u(x_{0})=\frac{-D(x_{0})+\sqrt{4A(x_{0})C(x_{0})+D(x_{0})^{2}}}{2A(x_{0})}$
$=\frac{aQ_{m}^{2k}x^{2-2k}-ax+x^{2}\sqrt{a^{2}+Q_{m}^{2k}x^{2-2k}+x^{2}-x}}{Q_{m}^{2k}x^{2-2k}+x^{2}-x}.$ (21)
In the above equation, $x_{0}$ represents the shortest distance that the light ray reaches the BH. Since we are mainly concerned with the behavior of light rays near the unstable photon orbit, that is, the case of $x_{0}\approx x_{m}$, the impact parameter corresponding to the unstable photon orbit can be expressed as $u(x_{m})$. The relationship between this impact parameter and the spin parameter $a$ and the hair parameter $Q_{m}$ can be visually presented in Figure 4. As the spin parameter $a$ or the short hair parameter $Q_{m}$ increases, the impact parameter $u(x_{m})$ gradually decreases, and this trend is consistent with the change trend of the photon-orbit radius. Here, we define the photon’s orbiting direction as counter clockwise. For the case of $a>0$ (the BH rotates counter clockwise), the photon’s orbiting direction is the same as the spin direction of the BH, and such an orbit is called a prograde orbit. Conversely, when $a<0$, the photon’s orbiting direction is opposite to the spin direction of the black hole, which is called a retrograde orbit. The radius of the photon’s prograde orbit is usually smaller than that of the retrograde orbit. This is because the spin effect enhances the stability of the photon closer to the BH on the prograde orbit through gravitational dragging. While on the retrograde orbit, the reverse dragging effect of the spin makes it more difficult for the photon to approach the BH, so the orbit radius is larger (see Figure 3).
Figure 4: The variation of the impact parameter with the spin parameter $a$ and the short hair parameter $Q_{m}$ in two BH models. From left to right, they correspond to the Kerr-Newman BH with $k=1$ and the rotating short-haired BH with $k=1.5$ respectively.
In the strong field limit, the deflection angle $\alpha_{D}$ can be given by Bozza (2002, 2003)
$\alpha_{D}(x_{0})=I(x_{0})-\pi,$ (22)
where
[…]
For the convenience of writing, it should be noted here that all capital letters in this paper represent functions of $x$, and all letters with the subscript “0” correspond to functions when $x=x_{0}$.
To handle the integral (23), we draw on the approaches presented in the references Bozza (2002); Tsukamoto (2017). Specifically, we expand the deflection angle in the vicinity of the unstable photon sphere. Additionally, we utilize an intermediate quantity re-defined by scholar Naoki Tsukamoto under the strong field limit to reconstruct the integral range. The form of this variable can be found in the definition given in Tsukamoto (2017), and its expression is
$z=1-\frac{x_{0}}{x}.$ (24)
This variable has been well applied in corresponding literature, such as Fu and Zhang (2022); Islam and Ghosh (2021). By using the above intermediate variable, equation (23) becomes
[…]
where
$R(z,x_{0})=\frac{2x^{2}}{x_{0}}\frac{\sqrt{B}(2A_{0}Au+A_{0}D)}{\sqrt{CA_{0}}\sqrt{4AC+D^{2}}},$ (26)
$f(z,x_{0})=\frac{1}{\sqrt{A_{0}-\frac{AC_{0}}{C}+\frac{u}{C}(AD_{0}-A_{0}D)}}.$ (27)
Evaluating the above expressions, we find that $R(z,x_{0})$ is positive definite everywhere from 0 to 1. However, $f(z,x_{0})$ diverges as $\lim_{z\rightarrow 0}f(z,x_{0})$. To avoid this issue, we can expand the denominator in a series up to the second-order term. Here, we define the expansion function as $K(z,x_{0})=A_{0}-\frac{AC_{0}}{C}+\frac{u}{C}(AD_{0}-A_{0}D)$, then equation (27) can be approximated as
$f_{0}(z,x_{0})=\frac{1}{\sqrt{\gamma_{0}(x_{0})+\gamma_{1}(x_{0})z+\gamma_{2}(x_{0})z^{2}}},$ (28)
where
$\gamma_{0}(x_{0})=K(z,x_{0})|_{z=0}=0,$ (29)
$\gamma_{1}(x_{0})=\left.\frac{\partial K(z,x_{0})}{\partial z}\right|_{z=0}=x_{0}\left(uD_{0}-C_{0}\right)\left(\frac{A_{0}^{\prime}C_{0}-A_{0}C_{0}^{\prime}}{C_{0}^{2}}\right)-uA_{0}x_{0}\left(\frac{D_{0}^{\prime}C_{0}-D_{0}C_{0}^{\prime}}{C_{0}^{2}}\right),$ (30)
$\gamma_{2}(x_{0})$ $=\frac{1}{2!}\left.\frac{\partial^{2}K(z,x_{0})}{\partial z^{2}}\right|_{z=0}=\frac{1}{2}\left[x_{0}^{2}\left(uD_{0}-C_{0}\right)\left(\frac{A_{0}^{\prime\prime}C_{0}-A_{0}C_{0}^{\prime\prime}C_{0}^{2}-2C_{0}C_{0}^{\prime}(A_{0}^{\prime}C_{0}-A_{0}C_{0}^{\prime})}{C_{0}^{4}}\right)\right.$
$\left.-uA_{0}x_{0}^{2}\left(\frac{D_{0}^{\prime\prime}C_{0}-D_{0}C_{0}^{\prime\prime}C_{0}^{2}-2C_{0}C_{0}^{\prime}(D_{0}^{\prime}C_{0}-D_{0}C_{0}^{\prime})}{C_{0}^{4}}\right)\right].$ (31)
In the above equations, the symbol denotes the first derivative, and ′′ denotes the second derivative.
According to the description of the strong - field limit, when light approaches the vicinity of the photon orbit radius $x_{m}$, the deflection angle of the light increases sharply. At this point, the deflection angle can be described by an analytical expansion as Bozza (2002, 2003); Tsukamoto (2017)
$\alpha_{D}(b)=-\bar{a}\log\left(\frac{\theta D_{OL}}{u_{m}}-1\right)+\bar{b}+O((b-b_{m})\log(b-b_{m})).$ (32)
The angular separation between the image and the lens can be approximated as $\theta\approx\frac{u}{D_{OL}}$, where $D_{OL}$ is the distance from the observer to the lens. The corresponding lens coefficients in equation (32) are
$\bar{a}=\frac{R(0,x_{m})}{2\sqrt{\gamma_{2}(x_{m})}},$ (33)
and
$\bar{b}=-\pi+I_{R}(x_{m})+\bar{a}\log\left(\frac{cx_{0}^{2}}{u_{m}^{2}}\right),$ (34)
where
[…]
and and $c$ is the expansion coefficient of the impact parameter at $u_{m}$:
$u-u_{m}=c(x_{0}-x_{m})^{2}.$ (36)
Theoretically, substituting the corresponding expressions should yield an analytical expression for the deflection angle. However, due to the excessive complexity and length of the expression, we have instead plotted the relationships among the lensing coefficients, the short hair parameter, and the spin parameter. As shown in Figure 5, the deflection coefficient $\bar{a}$ increases with the increase of the spin parameter $a$, and the presence of the hair parameter makes this growth more rapid. When the hair parameter vanishes, i.e., $Q_{m}=0$, the rotating short-haired BH degenerates into a Kerr BH. The behavior of the deflection coefficient $\bar{b}$ is exactly opposite to that of the coefficient $\bar{a}$, and the presence of hair causes the corresponding coefficient to decrease more rapidly. When the parameters $a=0$ and $Q_{m}=0$, the short-haired BH degenerates into a Schwarzschild BH. At this point, the lensing coefficient $\bar{a}=1$ and $\bar{b}=-0.4002$, which is in good agreement with the lensing coefficient values of the Schwarzschild BH Bozza (2002)(see the intersections of the red contour lines and the abscissa in the left and right plots of Figure 5, as well as Table 1 ). When the parameters $a=0$, $k=1$, and $Q_{m}\neq 0$, the short-haired BH degenerates into a Reissner-Nordström (RN) BH. Our calculated results are in excellent agreement with those calculated by Bozza Bozza (2002) (Table 1).
In addition, in Figure 6, we have plotted the variation of the deflection angle of the short-haired BH with the impact parameter under different values of the hair parameter and the spin parameter. Evidently, as the impact parameter $u$ continuously decreases, different hair parameters always correspond to different divergence points (the points on the dashed lines in Figure 6 correspond to the values of the impact parameter at divergence). Among them, the deflection angle of the Kerr BH is greater than that of the rotating short-haired BH under the same impact parameter (the black dashed line in Figure 6 represents the case of the Kerr BH). Overall, the presence of the hair parameter causes the deflection angle to diverge at a smaller impact parameter. Meanwhile, the variation trends of both the lensing coefficients $\bar{a}$ and $\bar{b}$, as well as the deflection angle, are similar to those of the standard Kerr BH.
Figure 5: The variation of the lensing coefficients $\bar{a}$ and $\bar{b}$ with the spin parameter $a$ and the hair parameter $Q_{m}$. Here, $k=1.5$ corresponds to the case of the rotating short-haired BH.
Lensing Coefficients
$a$ $Q_{m}$ $\bar{a}$ $\bar{b}$ $u_{m}/R_{s}$ $u_{m}/R_{s}$
$k=1$ $k=1.5$ $k=1$ $k=1.5$ $k=1$ $k=1.5$
-0.2 0 0.8796 0.8796 -0.2810 -0.2810 2.9788 2.9788
0.3 0.9048 0.8911 -0.2570 -0.2824 2.8331 2.9543
0.4 0.9354 0.9098 -0.2382 -0.2875 2.7043 2.9186
0.5 1.0152 0.9533 -0.2388 -0.3123 2.5054 2.8533
0 0 1.0000 1.$0$0.4002 -0.4002 2.5981 2.5981
0.3 1.0518 1.0220 -0.3965 -0.4124 2.4294 2.5658
0.4 1.1232 1.0609 -0.4136 -0.4408 2.2730 2.5176
0.5 1.4142 1.1714 -0.7332 -0.5677 2.0000 2.4231
0.2 0 1.2209 1.2209 -0.7700 -0.7700 2.1686 2.1686
0.3 1.3717 1.2802 -0.9488 -0.8508 1.9534 2.1200
0.4 1.6922 1.4096 -1.4988 -1.0636 1.7232 2.0413
0.5 - - - - - -
Table 1: The corresponding values of the lensing coefficients under specific spin parameters and hair parameters. When $a=0$ and $Q_{m}=0$, the BH degenerates into a Schwarzschild BH. When $a=0$ and $Q_{m}\neq 0$, the BH becomes a Reissner-Nordström (R-N) BH. When $a\neq 0$ and $Q_{m}=0$, the BH is a Kerr BH. The parameter $k=1$ corresponds to the Kerr-Newman BH, and $k=1.5$ corresponds to the rotating short-haired BH.
Figure 6: The variation of the deflection angle $\alpha_{D}$ of a rotating short-haired BH with the impact parameter $u$ under the strong deflection limit. From left to right, the spin parameters are $a=-0.2$ and $a=0.2$.
IV Observational Effects of Strong Gravitational Lensing
IV.1 Characteristic Observation Values and Time Delays in Strong Lensing Effects
In gravitational lensing, if the positions of the lens and the light source are known, the specific positions of the images can be calculated using the lens equation. The lens equation is given in Virbhadra and Ellis (2000), and later approximated using the small-angle approximation Bozza (2008b); Bozza et al. (2001)
$\beta=\theta-\frac{D_{LS}}{D_{OS}}\Delta\alpha_{n}.$ (37)
Here, $\beta$ represents the angle between the source and the lens axis, $\theta$ represents the angle between the image and the lens axis, $D_{LS}$ is the distance from the source to the lens, and $D_{OS}$ is the distance from the source to the observer. The relationship among the source, lens (black hole), and observer is given by $D_{OS}=D_{OL}+D_{LS}$. It is worth noting that, since the rotating short-hair black hole is asymptotically flat, for convenience of analysis, the source, lens, and observer can be set on the same line. In the above equation, $\Delta\alpha_{n}=\alpha(\theta)-2n\pi$ represents the remaining deflection angle after the light ray has looped around the black hole $n$ times.
To approximate the deflection $\Delta\alpha_{n}$, we use the analysis from Bozza (2003). For the $n$-th image, the relationship is given by
$\theta_{n}=\theta_{n}^{0}+\Delta\theta_{n},$ (38)
where
$\theta_{n}^{0}=\frac{u_{m}(1+e_{n})}{D_{OL}},$ (39)
$e_{n}=\exp\left(\frac{\bar{b}-2n\pi}{\bar{a}}\right),$ (40)
$\Delta\theta_{n}=\frac{D_{OS}}{D_{LS}}\frac{u_{m}e_{n}}{\bar{a}D_{OL}}(\beta-\theta_{n}^{0}).$ (41)
Here, $\theta_{n}^{0}$ corresponds to the angle $\theta$ when $\alpha=2n\pi$. By combining these relationships, the approximate position of the $n$-th image can be obtained as Bozza (2002)
$\theta_{n}=\theta_{n}^{0}+\frac{u_{m}e_{n}(\beta-\theta_{n}^{0})D_{OS}}{\bar{a}D_{LS}D_{OL}}.$ (42)
Note that the relativistic image obtained from equation (42) is only for the image on one side. For the image on the other side, we can obtain it in the same form by using $-\beta$. It can be seen from equation (42) that the latter term is just a correction to $\theta_{n}^{0}$.
Of course, in a lensing system, apart from the corresponding angular positions, the magnification is also an observable physical quantity of great significance. The magnification of the $n$th image can be expressed as Bozza (2002); Virbhadra et al. (1998); Virbhadra and Keeton (2008)
$\mu_{n}=\left(\frac{\beta}{\theta}\frac{d\beta}{d\theta}\right)^{-1}\Bigg{|}_{\theta_{n}^{0}}=\frac{u_{m}^{2}(1+e_{n})D_{OS}}{\bar{a}\beta D_{LS}D_{OL}^{2}}e_{n}.$ (43)
It is evident from the above expression that due to the presence of the $e_{n}$ term, the magnification of the image decreases exponentially with increasing $n$. Additionally, when $\beta$ is very small and approaches zero, we can obtain relatively bright images. In other words, images are easier to observe when $\beta\to 0$. Clearly, the magnification is largest when $n=1$, meaning the image is the brightest at this point. Therefore, if we resolve the first-order image $\theta_{1}$ (the dominant image) and represent the other unresolved images as $\theta_{\infty}$, we can obtain several interesting observables Bozza (2002)
$\theta_{\infty}=\frac{u_{m}}{D_{OL}},$ (44)
$S=\theta_{1}-\theta_{\infty}=\theta_{\infty}\exp\left(\frac{\bar{b}-2\pi}{\bar{a}}\right),$ (45)
$r_{\text{mag}}=\frac{\mu_{1}}{\sum_{n=2}^{\infty}\mu_{n}}\approx\frac{5\pi}{\bar{a}\ln(10)}.$ (46)
Here, $\theta_{\infty}$ represents the angular position of the unresolved bundled images, $s$ denotes the angular separation between the first resolved image (the outermost image) and the bundled images, and $r_{\text{mag}}$ is the brightness ratio between the outermost image and the remaining bundled images. A detailed discussion of these observables is presented in Section IV.2.
In addition to the observational effects mentioned above, time delay is also an important physical quantity. In the study of gravitational lensing effects, when light passes by a BH, its path bends due to gravity, forming multiple images. These images, because of different light paths, reach the observer at different times, which is known as time delay. This phenomenon can be used to determine the geometric scale and mass of the lensing system in observations and to estimate the Hubble constant in a cosmological context, such as Birrer et al. (2024); Treu et al. (2022); Grillo et al. (2018).Here, we use the method proposed by Bozza and Mancini for calculating time delay in the strong field limit Bozza and Mancini (2004). For a lensing system aligned on a straight line ($\beta=0$) and two relativistic images on the same side of the lens (the $p$-th and $q$-th images), the time delay formula is given by Bozza and Mancini (2004)
$\Delta T_{p,q}=2\pi(p-q)\frac{\tilde{a}}{\bar{a}}+2\sqrt{\frac{A_{m}u_{m}}{B_{m}}}\sqrt{u_{m}}\left(e^{-\frac{\bar{b}-2p\pi}{2\bar{a}}}-e^{-\frac{\bar{b}-2q\pi}{2\bar{a}}}\right).$ (47)
Analyzing the above equation, the first term represents the geometric time delay, which mainly depends on the number of loops the light makes around the black hole. The second term represents the time dilation effect of the light in the gravitational field. It is evident that the time delay is primarily determined by the difference in the number of loops the light makes around the BH, so the dominant term is the first term. Thus, the above equation can be approximated as
$\Delta T_{p,q}\approx 2\pi(p-q)\frac{\tilde{a}}{\bar{a}}=2\pi(p-q)u_{m}=2\pi(p-q)\theta_{\infty}D_{OL}.$ (48)
In the above discussion, we put forward several interesting observables, such as three key observable quantities and time delays. Based on these expressions, we can evaluate the observational results of the lensing effect in the real cosmic environment. In Section IV.2, we will take the known supermassive BHs M87* and Sgr A* in the real universe as our observation subjects. By using the relevant parameters of these actual BHs, we will assess the observational values of the lensing effect in the context of the rotating short-haired BH.
IV.2 Evaluating the Observability of Supermassive Black Holes
[…]
For the three interesting observables, when regarding the rotating short-haired BH as a candidate for M87* and Sgr A*, we carried out numerical calculations and presented the changes of these observables, as depicted in Figure 7. We selected the case of the rotating short-haired BH corresponding to the parameter $k=1.5$, and calculated the angular position $\theta_{\infty}$ of the relativistic image, the angular distance $s$, and the image magnification $r_{\text{mag}}$. The results show that, whether for Sgr A* or M87*, the variation trends of the observables are identical. Specifically, the angular position $\theta_{\infty}$ of the relativistic image decreases as the spin parameter $a$ and the hair parameter $Q_{m}$ increase. The angular distance $s$ increases with the increase of $a$ and $Q_{m}$. The image magnification $r_{\mathrm{mag}}$ gradually decreases as $a$ increases. In particular, as $Q_{m}$ increases, the variation trend becomes more significant (this is reflected by the density of the contour lines in the figure). By combining the data in Table 2, it can be seen that for M87*, the range of influence of the hair parameter $Q_{m}$ on the angular position of the relativistic image is $22.7113\ \mu as>\theta_{\infty}(M87*)>13.9162\ \mu as$. This range lies within the measurement range of $(42\pm 3)\ \mu as$ when the EHT observes the diameter of M87* Akiyama et al. (2019a). Regarding Sgr A*, the range of influence of the hair parameter is $29.3132\ \mu as>\theta_{\infty}(SgrA*)>17.9614\ \mu as$, and this range coincides with the measured value of $(51.8\pm 2.3)\ \mu as$ for the diameter of the Sgr A* shadow observed by the EHT Akiyama et al. (2022a). Obviously, these observed values are all at the $\mu as$ level, matching the observation accuracy of the existing EHT. Regrettably, though, the resolution of the current equipment is insufficient to detect these differences.
As can be seen from Table 3, under the same spin parameter conditions, for M87*, the deviations in observables between the rotating short-haired BH and the Kerr BH ($Q_{m}=0$) are $\delta\theta_{\infty}\approx 3\ \mu as$, $\delta S\approx 0.1272\ \mu as$, $\delta r_{mag}\approx 0.8980$; and the deviations between the rotating short-haired BH and the Kerr-Newman black hole ($k=1$) are $\delta\theta_{\infty}\approx 2.4249\ \mu as$, $\delta S\approx 0.0535\ \mu as$, $\delta S_{mag}\approx 1.1001$. For Sgr A*, the deviations between the rotating short-haired BH and the Kerr BH are $\delta\theta_{\infty}\approx 3.3786\ \mu as$, $\delta S\approx 0.1642\ \mu as$, $\delta r_{mag}\approx 0.8980$; and the deviations between the rotating short-haired BH and the Kerr-Newman BH are $\delta\theta_{\infty}\approx 3.1299\ \mu as$, $\delta S\approx 0.0690\ \mu as$, $\delta S_{mag}\approx 1.1001$. Although these deviations are all at the $\mu as$ level (especially the angular position), the angular resolution of the current EHT is approximately $20\ \mu as$, so it is impossible to distinguish these small differences Akiyama et al. (2019c). Therefore, to distinguish between the rotating short-haired BH and the Kerr BH or the Kerr-Newman BH may depend on the next-generation EHT. Once these differences can be distinguished in the future, it will provide us with further information about the short-hair parameter $Q_{m}$, and also offer an opportunity to test the no-hair theorem.
Figure 7: Variations of the Three Observables in M87* and Sgr A* with the Spin Parameter and Hair Parameter. The first row corresponds to M87*, and the second row corresponds to Sgr A*. Here, k = 1.5 represents the rotating short-haired BH.
$M87^{*}$ $SgrA^{*}$ $M87^{*}orSgrA^{*}$
$k=1$ $k=1.5$ $k=1$ $k=1.5$ $k=1$ $k=1.5$
$a$ $Q_{m}$ $\theta_{\infty}(\mu as)$ $S(\mu as)$ $\theta_{\infty}(\mu as)$ $S(\mu as)$ $\theta_{\infty}(\mu as)$ $S(\mu as)$ $\theta_{\infty}(\mu as)$ $S(\mu as)$ $r_{mag}$ $r_{mag}$
-0.2 0 22.7113 0.0130 22.7113 0.0130 29.3132 0.0168 29.3132 0.0168 7.7554 7.7554
0.3 21.6003 0.0157 22.5243 0.0142 27.8792 0.0202 29.0718 0.0183 7.5398 7.6559
0.4 20.6186 0.0193 22.2525 0.0163 26.6121 0.0250 28.7210 0.0210 7.2930 7.4980
0.5 19.1019 0.0310 21.7547 0.0215 24.6545 0.0400 28.0784 0.0278 6.7200 7.1564
0 0 19.8085 0.0248 19.8085 0.0248 25.5668 0.0320 25.5668 0.0320 6.8219 6.8219
0.3 18.5222 0.0323 19.5629 0.0279 23.9064 0.0417 25.2495 0.0361 6.4858 6.6750
0.4 17.3301 0.0446 19.1951 0.0339 22.3677 0.0576 24.7749 0.0438 6.0738 6.4300
0.5 15.2487 0.1068 18.4742 0.0533 19.6813 0.1378 23.8444 0.0688 4.8238 5.9239
0.2 0 16.5339 0.0512 16.5339 0.0512 21.3400 0.0661 21.3400 0.0661 5.5874 5.5874
0.3 14.8931 0.0764 16.1635 0.0614 19.2224 0.0986 20.8620 0.0793 4.9733 5.3286
0.4 13.1383 0.1322 15.5632 0.0848 16.9574 0.1707 20.0873 0.1095 4.0314 4.8396
0.5 - - 13.9162 0.1784 - - 17.9614 0.2303 - 2.2650
Table 2: Observed values corresponding to different hairy parameters $Q_{m}$ when considering M87* and Sgr A* as the rotating short-haired BH. Here, $k=1$ corresponds to the Kerr-Newman BH, and $k=1.5$ corresponds to the rotating short-haired BH.
$M87^{*}$ $SgrA^{*}$ $M87^{*}orSgrA^{*}$
$hair-Kerr$ $hair-KN$ $hair-Kerr$ $hair-KN$ $hair-Kerr$ $hair-KN$
$a$ $\delta Q_{m}$ $\delta\theta_{\infty}(\mu as)$ $\delta S(\mu as)$ $\delta\theta_{\infty}(\mu as)$ $\delta S(\mu as)$ $\delta\theta_{\infty}(\mu as)$ $\delta S(\mu as)$ $\delta\theta_{\infty}(\mu as)$ $\delta S(\mu as)$ $\delta r_{mag}$ $\delta r_{mag}$
-0.2 0.3 0.1870 0.0012 0.924 0.0015 0.2414 0.0015 1.1926 0.0019 0.0995 0.1161
0.4 0.4588 0.0032 1.6339 0.0030 0.5922 0.0041 2.1089 0.0040 0.2574 0.2050
0.5 0.9566 0.0085 2.6528 0.0095 1.2348 0.0109 3.4239 0.0122 0.5990 0.4364
0 0.3 0.2456 0.0031 1.0407 0.0044 0.3173 0.0041 1.3431 0.0056 0.1469 0.1892
0.4 0.6134 0.0092 1.865 0.0107 0.7919 0.0118 2.4072 0.0138 0.3919 0.3562
0.5 1.3343 0.0285 3.2255 0.0535 1.7224 0.0368 4.1631 0.0690 0.8980 1.1001
0.2 0.3 0.3704 0.0102 1.2704 0.0150 0.4780 0.0132 1.6396 0.0193 0.2588 0.3553
0.4 0.9707 0.0336 2.4249 0.0474 1.2527 0.0434 3.1299 0.0612 0.7478 0.8082
0.5 2.6177 0.1272 - - 3.3786 0.1642 - - 3.3224 -
Table 3: When considering M87* and Sgr A* as the rotating short-haired BH, the deviations in the observed values of the rotating short-haired BH from those in the cases of Kerr BH and Kerr-Newman BH. Here, $\delta(X)=X_{\mathrm{hair}}-X_{\mathrm{Kerr}}$ or $\delta(X)=X_{\mathrm{hair}}-X_{\mathrm{KN}}$.
$M87^{*}$ $SgrA^{*}$
$k=1$ $k=1.5$ $k=1$ $k=1.5$
$a$ $Q_{m}$ $\Delta T_{21}(h)$ $\Delta T_{21}(h)$ $\Delta T_{21}(min)$ $\Delta T_{21}(min)$
-0.2 0 332.1074 332.1074 12.2011 12.2011
0.3 315.8604 329.3723 11.6042 12.1006
0.4 301.5053 325.3982 11.0768 11.9546
0.5 279.3262 318.1182 10.2620 11.6872
0 0 289.6618 289.6618 10.6417 10.6417
0.3 270.8501 286.0676 9.9506 10.5097
0.4 253.4180 280.6904 9.3102 10.3121
0.5 222.9818 270.1475 8.1920 9.9248
0.2 0 241.7744 241.7744 8.8824 8.8824
0.3 217.7821 236.3585 8.0010 8.6834
0.4 192.1213 227.5810 7.0582 8.3610
0.5 - 203.4960 - 7.4761
Table 4: When considering M87* and Sgr A* as a rotating short-haired BH, the time delay $\Delta T_{21}$ between two images on the same side.
Figure 8: When considering M87* and Sgr A* as the rotating short-haired BH, the time delay deviation $\delta\Delta T_{21}$ of the same image between the rotating short-haired BH and the Kerr BH. Here, $\delta(X)=X_{\mathrm{hair}}-X_{\mathrm{Kerr}}$.
Figure 9: When M87* and Sgr A* are regarded as the rotating short-haired BH, the time delay deviation of the same image, $\delta\Delta T_{21}$, between the rotating short-haired BH and the Kerr BH. Here, $\delta(X)=X_{\mathrm{hairy}}-X_{K-N}$.
For the two relativistic images located on the same side of the BH, the time delay between the second relativistic image ($p=2$) and the first relativistic image ($q=1$) is presented in Table 4. The deviations between the rotating short-haired BH, the Kerr BH, and the Kerr-Newman BH are shown in Table 4 as well as Figures 8 and 9. Here, the time delay we consider is mainly determined by the optical path difference between relativistic images formed after light rays orbit the BH several times (see Figure 6. Obviously, near the divergence of the deflection angle, the light rays orbit more than $2\pi$). When considering M87* and SgrA* as a rotating short-haired BH, the time delay $\mathrm{\Delta}T_{2,1}$ between the second and the first relativistic images on the same side is such that the time delay of the former can be up to several hundred hours, while that of the latter can reach tens of minutes. Evidently, the time delay of the former is sufficient for astronomical observations, which provides a necessary condition for exploring the properties of the short-haired BH (see Table 4). Under the same circumstances, the time delay deviation amounts between the rotating short-haired BH on one hand, and the Kerr BH and the Kerr-Newman BH on the other hand: In the case of Sgr A* as the background, the time delay deviation between the rotating short-haired BH and the Kerr BH can reach up to more than two minutes (see Figure 8), while the time delay deviation between the rotating short-haired BH and the Kerr-Newman BH can reach up to more than four minutes (see Figure 9). In the case of M87* as the background, the time delay deviation between the rotating short-haired BH and the Kerr BH can reach up to more than fifty hours (see Figure 8), and the time delay deviation between the rotating short-haired BH and the Kerr-Newman BH can reach up to more than one hundred hours (see Figure 9). Overall, for Sgr A*, the time delay deviation is in the range of several minutes, which is difficult to be detected in astronomy. For M87*, the time delay can reach hundreds of hours. The time delay deviation between it (when considering M87* as a rotating short-haired BH) and the Kerr BH can reach over fifty hours, and the deviation between it and the Kerr-Newman BH can reach over one hundred hours. These time durations are sufficient for astronomical observations. Just from the perspective of time delay, if M87* is regarded as a candidate for a rotating short-haired BH, it is feasible to distinguish between traditional BHs and rotating short-haired BHs. However, the prerequisite for this is that the equipment used is capable of resolving the two relativistic images. Obviously, current equipment fails to meet this requirement. But the next generation EHT is expected to obtain such resolution. If these images can be distinguished in the near future, a lot of important physical information will be obtained, such as testing the no-hair theorem, exploring the properties of the hair parameter $Q_{m}$, and restricting the variation of parameters.
V Constraints from the EHT
Figure 10: After taking the rotating short-haired BH model as a candidate model for M87* and Sgr A*, the parameter space ($Q_{m}/M,a/M$) is constrained using the EHT data. The left hand figure corresponds to M87*, and the right hand figure corresponds to Sgr A*. The red dashed line represents the boundary of the corresponding first confidence interval. Here, the inclination angle $\theta_{o}=\frac{\pi}{2}$.
In the preceding section, we delved into the influence exerted by the hair parameter on the strong gravitational lensing phenomenon. The hair parameter acts by modifying the spacetime structure in the vicinity of the rotating short-haired BH, thereby inducing alterations in the propagation trajectories of light rays within the BH’s gravitational field. As light rays approach the event horizon, the intense gravitational field of the black hole causes the light rays to undergo deflection, giving rise to the gravitational lensing effect. This may potentially result in the formation of multiple images or regions of enhanced luminosity, as elaborated in detail in the previous section. With further deflection of the light rays, a black hole shadow is ultimately formed. The geometric properties of the shadow of a rotating short-haired BH are not only contingent upon the black hole’s spin but are also significantly influenced by the hair parameter. Consequently, while having an impact on the propagation paths of light rays, the hair parameter exerts a far-reaching influence on both the shape and size of the black hole shadow. A comprehensive discussion regarding the effect of the hair parameter on the shape of the shadow surrounding a rotating short-haired BH can be found in the reference Tang and Xu (2022). Therefore, our emphasis does not lie in a detailed exploration of the effects on the shadow shapes. Instead, we are committed to systematically constraining the relevant parameter space of the rotating short-haired BH by leveraging the shadow data of supermassive BHs obtained from the observations of the EHT.
In 2019, the EHT successfully captured the image of the black hole shadow of the supermassive BH M87* located at the center of the elliptical galaxy M87, which brought revolutionary progress to the observational and theoretical research of supermassive BHs. The EHT team measured the diameter of the black hole shadow of M87* to be $\theta_{d}=42\pm 3\ \mu\mathrm{as}$ Akiyama et al. (2019b, a, d). This result not only provides us with precise observational evidence of the geometric structure of the BH but also offers evidence for verifying the validity of general relativity in extreme gravitational fields. Subsequently, in 2022, the EHT team further enhanced the observational technology and data processing capabilities and successfully measured the diameter of the black hole shadow of the supermassive BH Sgr A* located at the center of the Milky Way. The result is $\theta_{d}=51.8\pm 2.3\ \mu\mathrm{as}$ Akiyama et al. (2022a). These data play a crucial role in the detection of gravitational theories and parameter constraints, such as Jusufi et al. (2022); Zahid et al. (2023); Afrin et al. (2023); Raza et al. (2024); Sánchez (2024); Zahid et al. (2023).
For axisymmetric black holes, the geometric shape of the black hole shadow is influenced by the spin parameter. To more accurately describe the shape of the shadow, an astronomical coordinate system Hioki and Maeda (2009) is usually introduced, which is defined as
$X=\lim_{r_{o}\rightarrow\infty}\left(-r_{o}^{2}\sin\theta_{o}\frac{d\phi}{dr}\right),$ (49)
$Y=\pm\lim_{r_{o}\rightarrow\infty}\left(r_{o}^{2}\frac{d\theta}{dr}\right),$ (50)
where $r_{o}$ and $\theta_{o}$ are the distance and inclination angle of the observer in the Boyer - Lindquist coordinate system, respectively. If considering the equatorial plane ($\theta=\frac{\pi}{2}$) and combining with the metric (2) of the rotating short-haired BH, then formulas (49) and (50) can be rewritten as
$X=\frac{4r\Delta-\Delta^{\prime}r^{2}+a^{2}}{a\Delta^{\prime}},$ (51)
$Y=\pm\frac{\sqrt{r^{2}(a^{2}-\Delta)(16\Delta-r^{2}\Delta^{\prime 2}+8r\Delta\Delta^{\prime})}}{a^{2}\Delta^{\prime 2}}.$ (52)
Here, the prime (’) indicates the derivative with respect to $r$, and $\Delta$ is a function term associated with the rotating short-haired BH. The detailed derivation process of this part can be found in corresponding literature, such as Tang and Xu (2022); Zahid et al. (2023); Sarikulov et al. (2022), etc.
With these preparations, we can model the rotating short-haired BH model as M87* or Sgr A*, so as to constrain the parameter space $(Q_{m}/M,a/M)$ of the rotating short-haired BH by using the data observed by the EHT. As shown in the left hand panel of Figure 10, considering the rotating short-haired BH model as a candidate for M87* and combining with the EHT observational data, we have constrained the parameter space of $Q_{m}$ (hair parameter) and $a$ (spin parameter) within the first confidence interval. The constrained region is enclosed by the red dashed line. The red curve in the figure, corresponding to an angular diameter $\theta_{d}=39\ \mu as$, represents the boundary of the first confidence interval. Within this confidence interval, the EHT data provides an effective constraint on the parameter space $(Q_{m}/M,a/M)$ of the rotating short-haired BH. Specifically, the constraint on the hair parameter is influenced by the spin parameter. As the $a$ increases, the constrained range of the $Q_{m}$ gradually shrinks.
Similarly, when applying the rotating short-haired BH model to Sgr A*, we can also utilize the EHT observational data of the Sgr A* shadow to constrain the parameter space of $Q_{m}$ and $a$ within the first confidence interval. The region enclosed by the red dashed line is the constrained interval. In the right figure of Figure 10, the red dashed line, corresponding to an angular diameter $\theta_{d}=49.5\ \mu as$, represents the boundary of the first confidence interval. As shown in the right figure, the constraints on the parameter space in Sgr A* are similar to those in M87*. Both the spin parameter $a$ and the hair parameter $Q_{m}$ are significantly constrained, and the constrained range of $Q_{m}$ gradually narrows as $a$ increases.
Overall, whether using the shadow data of M87* or Sgr A* to constrain the parameter space $(Q_{m}/M,a/M)$ of the rotating short-haired BH model, our analysis results show that the parameter space of the rotating short-haired BH model is highly consistent with the range of the observational data from the EHT. Therefore, through a systematic analysis of the EHT observational results of M87* and Sgr A* within the first confidence interval, we cannot rule out the rotating short-haired BH model as a potential candidate for real universe black holes. In addition, although the current observational accuracy still has limitations, the EHT data has provided a certain detectability for the rotating short-haired BH model. With the gradual improvement of observational techniques, future data may help refine the parameter space of this model and provide new clues for distinguishing different black hole models. This result supports the potential of the rotating short-haired BH model as a description of real universe black holes.
VI Discussion and conclusions
The gravitational lensing effect has become an important tool for studying the spacetime structure of black holes. In particular, the 2019 image of M87* revealed its high consistency with the Kerr metric, strongly supporting the theoretical expectation that black holes follow the Kerr metric. However, the existence of accretion flows, dark matter halos, and non-vacuum environments poses significant challenges to the isolated Kerr black hole model. As a natural extension of the Kerr solution, the rotating short-haired BH model may provide a new theoretical framework for studying real black holes in the universe and their complex environments.
This paper investigates the strong gravitational lensing effect in the rotating short-haired BH and conducts a constraint analysis on relevant parameters. We find that the hair parameter significantly influences the black hole shadow and the gravitational lensing effect. Specifically, as the intensity of the short hair increases, the event horizon radius of the rotating short-haired BH decreases, and the photon orbit radius and the impact parameter decrease with the increase of the black hole spin. Meanwhile, the photon orbit radius and the impact parameter of the Kerr BH are larger than those of the rotating short-haired BHe, and the degree of deviation increases with the increase of the hair parameter. Regarding the lensing coefficients $\bar{a}$ and $\bar{b}$, the coefficient $\bar{a}$ shows a monotonically increasing trend, while the coefficient $\bar{b}$ shows a monotonically decreasing trend, and the hair parameter exacerbates these changes.
We further analyzed the observational manifestations of the rotating short-haired BH in M87* and Sgr A*. Whether for M87* or Sgr A*, the changing trends of these observables are consistent: the angular position of the relativistic image $\theta_{\infty}$ decreases as the spin parameter increases, the angular distance $s$ increases with the increase of the spin parameter, and the magnification of the relativistic image $r_{mag}$ decreases as the spin parameter increases. The presence of hair makes both $\theta_{\infty}$ and $s$ smaller than those in the case of a Kerr BH, and $r_{mag}$ is larger than that of a Kerr BH. The larger the hair parameter, the greater the degree of deviation. By analyzing the data in Table 2, it can be concluded that the hair parameter $Q_{m}$ has a remarkable impact on the angular position of the relativistic image. Its variation ranges are $22.7113\ \mu as>\theta_{\infty}(M87^{*})>13.9162\ \mu as$ and $29.3132\ \mu as>\theta_{\infty}(SgrA^{*})>17.9614\ \mu as$ respectively. These ranges are comparable to the observational results of the shadow diameters of M87* and Sgr A* obtained by the EHT Akiyama et al. (2019a, 2022b) (see Figure 7 and Table 2). Through comparing the rotating short-haired BH with the Kerr BH ($Q_{m}=0$) and the Kerr-Newman BH ($k=1$), we’ve found that the angular position deviation between the rotating short-haired BH and the Kerr BH can reach $\delta\theta_{\infty}\approx 3\ \mu as$ in the case of M87*, and $\delta\theta_{\infty}\approx 3.3786\ \mu as$ for Sgr A*. The deviations from the Kerr-Newman BH are $\delta\theta_{\infty}\approx 2.4249\ \mu as$ (for M87*) and $\delta\theta_{\infty}\approx 3.1299\ \mu as$ (for Sgr A*), respectively.
In addition, the time delay effect of the rotating short-haired BH also shows significant differences from other models. When simulating M87*, the time deviation between the rotating short-haired BH and the Kerr BH can reach more than 50 hours, while the deviation from the Kerr-Newman BH can reach more than 100 hours. For Sgr A*, the time deviations are 2 minutes and 4 minutes respectively. Apparently, if the rotating short-haired BH is taken as the M87* model, the time delay effect is sufficient to distinguish it from the Kerr BH or the Kerr-Newman BH, providing theoretical support for further exploration of the properties of the hair parameter. It is worth mentioning that the maximum value of the time delay deviation does not increase monotonically with the spin parameter, but reaches its maximum at a specific spin parameter (see Figures 8 and 9 and Table 4).
Finally, by integrating the observational data from the EHT, we systematically constrained the hair parameter of the rotating short-haired BH. The results indicate that, within the first confidence interval, the parameter space $(Q_{m}/M,a/M)$ of the rotating short-haired BH is highly consistent with the observational data (Figure 10). This suggests that the rotating short-haired BH cannot be excluded as a possible model for black holes in the real universe, providing theoretical support for further differentiating between different black hole models.
In summary, the hair parameter has a significant impact on the strong gravitational lensing effect and the black hole shadow. These impacts may lead to certain observational differences for distinguishing between different black hole models. Future research could further explore the influence of the hair parameter on the weak gravitational lensing effect and further verify the applicability of this model through high precision observations.
VII acknowledgements
We acknowledge the anonymous referee for a constructive report that has significantly improved this paper. This work was supported by the Special Natural Science Fund of Guizhou University (Grant No.X2022133), the National Natural Science Foundation of China (Grant No. 12365008) and the Guizhou Provincial Basic Research Program (Natural Science) (Grant No. QianKeHeJiChu-ZK[2024]YiBan027) .